Rank the following temperatures from warmest to coldest: (a) 0°C ; (b) 0°F ; (c) 0 K ; and (d) 77 K.
Answer #1

a)

T=0oC

b)

Relationship Between oC and oF

oC = (oF-32)*(5/9)

T=(0-32)*(5/9)= -17.78oC

c)

Relationship between K and oC is

oC=K- 273.15

T=0-273.15 =-273.15oC

d)

T=77-273.15 =-196.15oC

Therefore ranking is A>B>D>C